# Christian Peukert [^1] Corresponding author. Florian Abeillon Jérémie Haese Franziska Kaiser Alexander Staub

### v1

@@ Abstract
Human-created works represent critical data inputs to artificial intelligence (AI). Strategic behavior can play a major role for AI training datasets, be it in limiting access to existing works or in deciding which types of new works to create or whether to create new works at all. We examine creators’ behavioral change when their works become training data for AI. Specifically, we focus on contributors on Unsplash, a popular stock image platform with about 6 million high-quality photos and illustrations. In the summer of 2020, Unsplash launched an AI research program by releasing a dataset of 25,000 images for commercial use. We study contributors’ reactions, comparing contributors whose works were included in this dataset to contributors whose works were not included. Our results suggest that treated contributors left the platform at a higher-than-usual rate and substantially slowed down the rate of new uploads. Professional and more successful photographers react stronger than amateurs and less successful photographers. We also show that affected users changed the variety and novelty of contributions to the platform, with long-run implications for the stock of works potentially available for AI training. Taken together, our findings highlight the trade-off between interests of rightsholders and promoting innovation at the technological frontier. We discuss implications for copyright and AI policy.
@@ 1 Introduction
Data is a key input to artificial intelligence (AI) that enables societal and economic progress through the discovery of knowledge, enhancement of productivity, and expansion into new or existing markets (Wu et al., 2020; Brynjolfsson et al., 2023; Beraja et al., 2023; McElheran et al., 2024). The rapid proliferation of generative AI (genAI) technologies has enabled hundreds of millions of users to generate high-quality text (including software code), images, audio, and video at a negligible cost. All this is made possible by the vast stock of data available on the public internet, which together with computing power, is responsible for the tremendous performance improvements in recent large language models (Ho et al., 2024; Metz et al., 2024). However, access to a continuous flow of data remains crucial for the performance of a wide range of AI applications in dynamic social environments (Hadsell et al., 2020; He et al., 2011; Peukert et al., 2023; Valavi et al., 2022; Whang et al., 2023). Input data for genAI models, in particular, comes mostly from online content in the form of text, images, sound, and video produced by humans. Hence, human behavior and incentives for individuals and organizations to contribute to training datasets play a major role for the quality of genAI (Bauer et al., 2016; Sokol and Van Alstyne, 2021; Burtch et al., 2024; del Rio-Chanona et al., 2023; Quinn and Gutt, 2023).
@@
Questions of whether, how, and which data can or should be utilized to train AI have moved to the center of the policy debate. Regulatory frameworks – or the lack thereof – can significantly impact both the demand for and supply of data. With respect to input data, AI policy intersects with privacy law and competition policy (Jones and Tonetti, 2020; Farboodi et al., 2019). More recently, however, intellectual property law has emerged as a critical area of the policy discourse because input data may be subject to copyright protection (Eshraghian, 2020; Samuelson, 2023). While some jurisdictions allow exemptions for research and development in copyright law, there is general legal uncertainty (Fiil-Flynn et al., 2022; Henderson et al., 2023). Several high-profile lawsuits against AI developers allege, among other claims, direct copyright infringement by creating unauthorized copies of their works and using such copies as training data. [^2] See articles in the New York Times (http://tiny.cc/pxljxz) and Reuters (http://tiny.cc/txljxz). As a result, policymakers around the world are tasked with balancing innovation in AI and the interests of rightsholders. So far, however, there is not sufficient empirical evidence to guide policy (Peukert and Windisch, 2024).
@@
In this paper, we shine a light on how strategic behavior shapes the flow of data, studying the response of individual creators to their works being made available for the training of commercial AI. Our empirical setting is one of the largest stock photography websites in the world, Unsplash. In the summer of 2020, Unsplash released metadata of a subset of 25,000 photos explicitly for the training of commercial AI applications (hereafter the “LITE” dataset). This natural experiment allows us to identify the causal effects of being included in AI training data and characterize strategic responses of those affected. We compare the upload behavior of contributors whose works were included in this dataset to contributors whose works were not included. In addition, we investigate changes in the variety and novelty of user contributions due to the release of the LITE dataset. We measure variety and novelty by calculating the similarity of each new upload to all existing images one year prior to the treatment based on their associated keywords using natural language processing methods. This allows us to track how variety and novelty change over time for both treated and control users, before and after the release of the LITE dataset.
We find that treated users left the platform at a higher-than-usual rate. Conditional on remaining active, treated users substantially slowed down the rate of new uploads by about 40% per month.
@@
Our content analysis shows that within-users, uploads decrease in variety but not in novelty compared to the existing stock of images. Across users, however, the variety of uploaded images decreased by about 5% compared to the existing stock and uploaded images were about 30% less novel. This implies changes to the aggregated training dataset stemming mostly from changes in user composition rather than from a shift in individual behavior. We show that users with professional gear reduce their contributions significantly more, as do more successful users. However, our results do not imply that photographers leave the market entirely. We show that users’ behavioral changes only occur on Unsplash and are not mirrored in their activity on Instagram, the most popular image-sharing platform. Finally, we provide some evidence that the behavioral response is stronger, the more AI may be perceived as an economic threat. First, we show that the effect intensifies later when more capable genAI models become widely known to the public. Second, users who have multiple images in the training dataset reduce their activity almost twice as much compared to users who had only one image included.
Overall, we document that strategic behavior affects the size as well as aspects of the quality of the flow of data. A back-of-the-envelope calculation suggests that making the entire catalog available for commercial AI research (which is similar to a policy that would allow fair use of any copyrighted material) would have reduced the flow of data by half. At the same time, the flow would become increasingly similar to the stock, with the number of very similar images tripling within a year.
@@
Our contribution is twofold. First, we provide the first large-scale empirical evidence on strategic behavior in the supply of data. This adds to a growing literature that discusses the effects of AI on workers and platform contributors (Acemoglu et al., 2022; Doshi and Hauser, 2023; Eloundou et al., 2023; Felten et al., 2023; Huang et al., 2023; Quinn and Gutt, 2023). In particular, we highlight upstream consequences of the proliferation of genAI models, which is an under-explored topic (Lin, 2024). Second, our setting allows us to estimate the causal effects of adding human-made works to training datasets that can be used for the development of commercial AI applications. Thereby we are able to provide important empirical evidence to a primarily theoretical debate on policy-making in the age of AI (Fiil-Flynn et al., 2022; Samuelson, 2023; Henderson et al., 2023; Gans, 2024; Yang and Zhang, 2024; Wang et al., 2024).
@@ 2 Background and related literature
Regulation can affect the demand for data by defining rules about which data can or should be used for AI applications. For example, privacy law and competition policy can reduce firms’ abilities to track consumer behavior (Johnson, 2022), require firms to elicit consent from consumers (Godinho de Matos and Adjerid, 2022) or share data with competitors (Prüfer and Schottmüller, 2021; Martens et al., 2024). Access to data has direct consequences for the output of AI applications. The value of data, for example for prediction exercises, varies with its quality and quantity (Lei et al., 2023; Neumann et al., 2019; Peukert et al., 2023; Sun et al., 2023).
@@
Concretely, the new AI Act of the European Union will demand that “Training, validation and testing datasets shall be relevant, sufficiently representative, and to the best extent possible, free of errors and complete in view of the intended purpose. They shall have the appropriate statistical properties, including, where applicable, as regards the persons or groups of persons in relation to whom the high-risk AI system is intended to be used”. [^3] See Article 10 of the Proposal of the European Parliament and of the Council on the Artificial Intelligence Act from January 2024, https://data.consilium.europa.eu/doc/document/ST-5662-2024-INIT/en/pdf. However, such provisions also have intersections and perhaps even contradictions with intellectual property law. In the EU, for example, although copyright law was only recently modernized with the Directive on Copyright in the Digital Single Market in 2019, discussions about further reform are in full swing as part of legislation on the regulation of AI. When rightsholders can exercise exclusive rights and restrict access to certain source materials, copyright may, in turn, create or promote biased AI systems (Levendowski, 2018).
@@
On the other hand, the supply of data is discussed less. With the abundance of data online, researchers and AI developers can potentially access a large stock of data, most of which was created before the dawn of AI. In general, surprisingly little is known about input datasets for AI applications. Interesting empirical work comes from the computer science community. Birhane and Prabhu (2021) study a widely-used training dataset of images concerning privacy issues (the presence of human faces), sexually explicit content, and measures of the age and gender distribution of depicted humans. In related recent work, Guilbeault et al. (2024) study gender bias in about one million images downloaded from the web. In many applications, access to a large stock of historical data is sufficient. Yet, in dynamic settings, where the stock of data can become outdated, a consistent flow of data is key for algorithmic performance (Hadsell et al., 2020; He et al., 2011; Valavi et al., 2022; Peukert et al., 2023; Whang et al., 2023). However, the flow of data is endogenously determined by strategic behavior, which is, in turn, influenced by institutional settings such as privacy law, competition policy, and intellectual property law.
@@
Key criteria that define the quality of training datasets concern the variety and novelty of information in the dataset. To illustrate this, consider two examples. First, imagine a panel dataset that an econometrician would analyze. We can increase variety by adding more cross-sections, e.g. more firms, and we can increase novelty by expanding the time series, i.e. adding more years per firm. As a second example, consider a dataset with photos of New York City. We can add to the variety dimension by adding images of buildings from every neighborhood, from different angles, at different times of the day, etc. However, as some buildings make way for new developments, we need to add novel images of the same concept, e.g. the skyline, to ensure a continued high-quality dataset. Variety and novelty can be measured by changes coming in the flow of incoming data points and their relation to the existing stock of data. As a result, strategic behavior that alters the flow of data can also alter key dimensions of dataset quality and, therefore, the output quality of AI applications.
@@
Recent theoretical work is useful as a starting point to think about policy options to solve the trade-off between copyright protection and AI innovation (Gans, 2024; Yang and Zhang, 2024). Digital goods, such as user-generated content or creative works, tend to have high fixed costs and near-zero costs of copying. Intellectual property rights or alternative business models create rents that allow creators to recoup their fixed-cost investments. The proliferation of genAI systems suggests two potential outcomes: AI might reduce these monopoly rents, leading humans to produce only works with lower fixed costs, or AI might increase human productivity, effectively reducing fixed costs and requiring lower monopoly rents to incentivize creation. These effects can depend on factors such as the bargaining power of rights holders (Gans, 2024) and the existing body of work in a particular domain (Yang and Zhang, 2024). A potential solution is a licensing mechanism that compensates rightsholders with a royalty scheme that is determined by Shapley values of pieces of training data that end up in the final outcome (Wang et al., 2024).
@@
Strategic behavior of contributors to training datasets, especially when large parts of the public internet can be considered as a huge training dataset, can have intricate long-run effects on online communities. Evidence from platforms like Stack Overflow suggests that the introduction of genAI systems like ChatGPT can lead to reduced community participation, a decrease in the number of questions and users as well as a decline in the quantity and quality of answers (Burtch et al., 2024; Quinn and Gutt, 2023; Gallea, 2023), and an overall decrease in activity (del Rio-Chanona et al., 2023). Three recent working papers are most directly related to our paper. Quinn and Gutt (2023) discuss how the quality of training data can change because of changes in user behavior. In the context of Stack Overflow, they show that changed user behavior affects the underlying quality of the training data for the next iterations of AI models. Huang et al. (2023) investigate the impact of genAI on content creators across two major Asian platforms, Lofter and Graffiti Kingdom, through quasi-experiments. The study reveals that introducing genAI tools on Lofter decreased creator activity, whereas banning these tools on Graffiti Kingdom had a mixed impact: Non-churning creators increased their activity, while that of individuals with copyright concerns decreased. Similarly, Lin (2024) studies the online art platform DeviantArt and shows that contributors decreased their publication volume by 22% after the platform introduced a genAI feature that let users generate art with a click of a button. In summary, while evidence on the impact of genAI trained on vast and intransparent datasets is growing, the lack of research on contributors’ responses to the explicit use of their work for the development of commercial AI models forms the backdrop of our study.
3 Empirical setting
3.1 Unsplash
Unsplash is a stock photo platform hosting approximately 6 million photos and illustrations, submitted by roughly 360,000 contributors. According to Similarweb, Unsplash is among the top 5 most visited websites in the “photography” category globally, receiving about 30 million visits each month. Over the past decade, Unsplash has recorded a cumulative total of 1 trillion image views, coming from tens of millions of individual users and thousands of API partners.
Unlike other stock photography websites, the vast majority of images on Unsplash are made available under a permissive license, which allows all images on the platform to be used freely for both non-commercial and commercial purposes. Users can, of course, revoke this license by deleting individual images or their entire accounts at any time. Unlike other websites that make images available under similar licenses (e.g. Wikimedia), contributions on Unsplash are of a high quality. Partially, this is enabled by rigorous moderation efforts of an editorial team that hand-selects images that are prominently shown on the website. As of May 2023, the curation team has selected about 300,000 images for curation, which represents about 6% of all images available on the platform.
In October 2022, Unsplash launched a paid subscription service called Unsplash+. Among other things, after being accepted into the program as a contributor, users can receive monetary compensation for uses of the works. [^4] See https://unsplash.com/blog/contribute-to-unsplash/. Further, the Unplash+ license explicitly disallows the use of images for machine learning, AI, and biometric tracking technology. [^5] See https://unsplash.com/plus/license.

### v2

# Christian Peukert Florian Abeillon Jérémie Haese Franziska Kaiser Alexander Staub

@@ Abstract
Artificial intelligence (AI) systems rely heavily on human-generated data, yet the people behind that data are often overlooked. Human behavior can play a major role in AI training datasets, be it in limiting access to existing works or in deciding which types of new works to create or whether to create any at all. We examine creators’ behavioral change when their works become training data for commercial AI. Specifically, we focus on contributors on Unsplash, a popular stock image platform with about 6 million high-quality photos and illustrations. In the summer of 2020, Unsplash launched a research program and released a dataset of 25,000 images for commercial AI use. We study contributors’ reactions, comparing contributors whose works were included in this dataset to contributors whose works were not. Our results suggest that treated contributors left the platform at a higher-than-usual rate and substantially slowed down the rate of new uploads. Professional photographers and more heavily affected users had a stronger reaction than amateurs and less affected users. We also show that affected users changed the variety and novelty of contributions to the platform, which can potentially lead to lower-quality AI outputs in the long run. Our findings highlight a critical trade-off: the drive to expand AI capabilities versus the incentives of those producing training data. We conclude with policy proposals, including dynamic compensation schemes and structured data markets, to realign incentives at the data frontier.
@@ 1 Introduction
Data is an essential input to artificial intelligence (AI) which enables societal and economic progress through the discovery of knowledge, enhancement of productivity, and expansion into new or existing markets (Wu et al., 2020; Beraja et al., 2023; McElheran et al., 2024; Brynjolfsson et al., 2025). The rapid proliferation of generative AI (genAI) technologies has enabled hundreds of millions of users to generate high-quality text (including software code), images, audio, and video at a negligible cost. All this is made possible by the vast stock of data publicly available on the internet, which, together with computing power, is responsible for the tremendous performance improvements in recent large language models (Ho et al., 2024). However, access to a continuous flow of data remains crucial for performing a wide range of AI applications in dynamic social environments (He et al., 2011; Valavi et al., 2022a; Peukert et al., 2024). Input data for genAI models is primarily composed of online content produced by humans, in the form of text, images, sound, and video. Hence, the incentives that drive contributions to training datasets, coupled with institutional frameworks, play a major role for the quality of genAI (del Rio-Chanona et al., 2023; Burtch et al., 2024; Jia and Nagaraj, 2025; Quinn and Gutt, 2025).
@@
Questions of whether, how, and which data can or should be used to train AI models have moved to the center of the policy debate. Regulatory frameworks – or the lack thereof – can significantly impact both the supply of and demand for data. Concerning input data, AI policy intersects with privacy law and competition policy (Farboodi et al., 2019; Jones and Tonetti, 2020; Azoulay et al., 2024). More recently, however, intellectual property law has emerged as a critical area of the policy discourse because input data may be subject to copyright protection (Samuelson, 2023). While some jurisdictions allow exemptions for research and development in copyright law, general legal uncertainty remains (Fiil-Flynn et al., 2022; Henderson et al., 2023). Several high-profile lawsuits against AI developers allege, among other claims, direct copyright infringement by creating unauthorized copies of their works and using such copies as training data. [^1] See articles in the New York Times (http://tiny.cc/pxljxz) and Reuters (http://tiny.cc/txljxz). As a result, policymakers everywhere are tasked with balancing innovation in AI with the interests of rights holders (Appel et al., 2023; de Rassenfosse et al., 2025).
@@
In this paper, we highlight how human behavior can shape the flow of data, studying the response of users to their works being made available for the training of commercial AI applications. The empirical setting for our study is Unsplash, one of the largest stock photography websites. In the summer of 2020, Unsplash released metadata on, and enabled the bulk download of, a subset of 25,000 images explicitly to train commercial AI applications (hereafter the “LITE” dataset). This natural experiment allows us to identify the causal effects of being included in AI training data and characterize the responses of those affected. We compare the upload behavior of contributors whose works were included in this dataset to contributors whose works were not. We then explore the underlying mechanisms by studying heterogeneity across contributors. Finally, we investigate changes in the variety and novelty of contributions in response to the release of the LITE dataset. We measure variety and novelty by calculating the similarity of each new upload to all existing images one year prior to the treatment based on textual descriptions of their content using natural language processing (NLP) methods. This allows us to track how variety and novelty change over time for both treated and untreated users, before and after the release of the LITE dataset.
@@
We find that treated users left the platform at a higher-than-usual rate and, conditional on remaining active, substantially slowed down the rate of new uploads by about 40% per month. Our analysis of mechanisms suggests that this reduction is driven by the economic interests of contributors. We show that professional photographers and prolific users reduce their contributions significantly more. In addition, the behavioral response appears to be stronger when AI is perceived as an economic threat. Users with multiple images in the training dataset reduce their activity almost twice as much compared to users who only had one image included. Second, the effect intensifies around two years later when capable genAI models become widely known to the public. We also provide evidence that users shift away from the type of photography that is most prominently featured in the LITE dataset. However, our results do not suggest that photographers leave their profession altogether. For a subsample, we can show that users’ behavioral changes occur only on Unsplash and are not mirrored in their activity on Instagram, the most popular image-sharing platform.
Concerning potential long-run effects for AI training data, our results illustrate that user behavior affects the size and the quality of the flow of data. Our content analysis shows that within-users, uploads decrease in variety but not in novelty compared to the existing stock of images. Across users, the variety of uploaded images decreased by about 5% compared to the stock. Further, we find that affected users upload images that are about 30% less novel. This shows that changes to the aggregated training dataset stem primarily from changes in user composition but also from a shift in individual behavior. Extrapolating the effects to the aggregate suggests that making the entire catalog available for commercial AI research (which is similar to a policy that would exempt data mining from copyright law) would have reduced the supply of data by half. At the same time, the data flow would become increasingly similar to the stock, with the number of very similar images tripling within a year. Relating our findings to experimental research in computer science, we conclude that our estimates of a decrease in the quantity, variety, and novelty of the flow of training data translate into a non-negligible decrease in AI output quality.
We discuss managerial implications for platforms that face trade-offs between making data available for AI training and potential user responses, including measures to counteract decreased user activity. Furthermore, we outline policy proposals to address these negative consequences, including a dynamic compensation scheme for rights holders and data markets modeled on those used in targeted online advertising.
@@
The contributions of this research are twofold. First, we provide a large-scale empirical study of user behavior in the context of AI training data. Thereby, we add to a discussion of how changes in the upstream supply of human-generated data can affect genAI model outputs, which remains an under-explored research topic (Jia and Nagaraj, 2025). This supply-side perspective expands research on platform contributors’ role in the AI economy (del Rio-Chanona et al., 2023; Burtch et al., 2024; Yiu et al., 2024; Quinn and Gutt, 2025), but also complements research on the demand-side, discussing the downstream effects of genAI on creative tasks (Boussioux et al., 2024; Doshi and Hauser, 2024; Zhou and Lee, 2024), and the effects of AI on labor markets in general (Felten et al., 2023; Eloundou et al., 2024; Hartley et al., 2024; Brynjolfsson et al., 2025; Demirci et al., 2025). Second, our setting allows us to estimate the causal effects of adding human-made works to training datasets that can be used to develop commercial AI applications. Thereby, we provide empirical evidence to a primarily theoretical debate on policy-making in the age of AI (Fiil-Flynn et al., 2022; Henderson et al., 2023; Samuelson, 2023; Wang et al., 2024; Yang and Zhang, 2024; Gans, 2025).
@@ 2 Background and related literature
Regulation can affect the demand for data by defining rules about which data can or should be used for AI applications. For example, privacy law and competition policy can reduce firms’ abilities to track consumer behavior (Johnson, 2024), require firms to elicit consent from consumers (Godinho de Matos and Adjerid, 2022) or share data with competitors (Prüfer and Schottmüller, 2021; Lei et al., 2023). Access to data has a direct impact on AI applications, causing the value of data – especially for prediction tasks – to depend on both its quality and quantity (Neumann et al., 2019; Lei et al., 2023; Peukert et al., 2024; Sun et al., 2024).
@@
Concretely, the AI Act of the European Union (EU) demands that “Training, validation and testing datasets shall be relevant, sufficiently representative, and to the best extent possible, free of errors and complete in view of the intended purpose.” [^2] See Article 10(3) of the European Union Artificial Intelligence Act signed in June 2024, https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=OJ:L_202401689#art_10. However, such provisions also have intersections and perhaps even contradictions with intellectual property law. In the EU, for example, although copyright law was only recently modernized with the Directive on Copyright in the Digital Single Market in 2019, discussions about further reform are in full swing. When rights holders can exercise exclusive rights and restrict access to certain source materials, copyright may, in turn, create or promote biased AI systems (Levendowski, 2018).
@@
With the abundance of data online, researchers and AI developers can potentially access a large stock of data, most of which was created before the dawn of AI. In general, surprisingly little is known about input datasets for AI applications. Recent empirical work demonstrates that AI models perform better when trained on larger and more diverse datasets, including potentially unauthorized sources. Jia and Nagaraj (2025) show that large language models exhibit significantly improved performance on books included in the Books3 dataset – a collection of pirated books – compared to those not present in Books3. Their findings highlight how restricted data access due to copyright laws could impact the development and accuracy of AI models. Other empirical contributions come from the computer science community. Birhane and Prabhu (2021) study a widely-used image training dataset concerning privacy issues such as the presence of human faces, sexually explicit content, and measures of the age and gender distribution of depicted humans. Relatedly, Guilbeault et al. (2024) explore gender bias in about one million images downloaded from the web.
In many applications, access to a large stock of historical data is sufficient. Yet, in dynamic settings, where the stock of data can become outdated, a consistent flow of data is key for algorithmic performance (He et al., 2011; Valavi et al., 2022a; Peukert et al., 2024). However, the flow of data is endogenously determined by human behavior, which is, in turn, influenced by institutional settings such as privacy law, competition policy, and intellectual property law. This dynamic dimension is severely understudied.
@@
Key criteria that define the quality of training datasets, be it in terms of stock or flow, concern the variety and novelty of information in the dataset. To illustrate this, consider two examples. First, imagine a panel dataset that an econometrician would analyze. We can increase variety by adding more cross-sections, e.g., more firms, and we can increase novelty by expanding the time series, i.e., adding more years per firm. As a second example, consider a dataset with photos of New York City. We can add to the variety dimension by adding images of buildings from every neighborhood, from different angles and times of the day, etc. However, as some buildings make way for new developments, it becomes necessary to add novel images of the same concept, e.g., the skyline, to ensure a continued high-quality dataset. Thus, variety and novelty can be measured by changes in the flow of incoming data points and their relation to the existing stock of data. As a result, user behavior that impact the flow of data can also alter key dimensions of dataset quality and, therefore, the output quality of AI applications.
@@
Recent theoretical work is a useful starting point to think about policy options to solve the trade-off between creator interests and AI innovation (Yang and Zhang, 2024; Gans, 2025). Digital goods, such as user-generated content or creative works, tend to have high fixed costs and near-zero costs of copying. Intellectual property rights or alternative business models create rents that allow creators to recoup their fixed-cost investments. The proliferation of genAI systems suggests two potential outcomes: AI might reduce these monopoly rents, leading humans to produce only works with lower fixed costs, or AI might increase human productivity, effectively reducing fixed costs and requiring lower monopoly rents to incentivize creation. These effects can depend on factors such as the bargaining power of rights holders (Gans, 2025) and the existing body of work in a particular domain (Yang and Zhang, 2024). A potential solution is a licensing mechanism that compensates rights holders with a royalty scheme that is determined by Shapley values of pieces of training data that end up in the final outcome (Wang et al., 2024).
@@
The behavior of contributors to training datasets, especially when large parts of the public internet are considered to constitute a training dataset, can have long-run effects on online communities. Evidence from platforms like Stack Overflow suggests that the introduction of genAI systems like ChatGPT can lead to reduced community participation, a decrease in the number of questions and users as well as a decline in the quantity and quality of answers (Burtch et al., 2024), and an overall decrease in activity (del Rio-Chanona et al., 2023). Three recent working papers are most directly related to our paper. Quinn and Gutt (2025) discuss how the quality of training data can change because of the reactions of users to the introduction of genAI. In the context of Stack Overflow, they show that while users reduce their contributions, the underlying quality of the contributions, and thus future training data, improves which has effects for the next iterations of AI models. Huang et al. (2023) investigate the impact of genAI on content creators across two major Asian platforms, Lofter and Graffiti Kingdom, through quasi-experiments. The study reveals that introducing genAI tools on Lofter decreased creator activity, whereas banning these tools on Graffiti Kingdom had a mixed impact: The activity of non-churning creators increased, while that of contributors with copyright concerns decreased. Similarly, Lin (2024) studies the online art platform DeviantArt and shows that contributors decreased their publication volume by 22% after the platform introduced a genAI feature that let users generate art with a click of a button.
In summary, while evidence on the impact of genAI trained on vast and intransparent datasets is growing, the lack of research on contributors’ responses to the explicit use of their work for the development of commercial AI models forms the backdrop of our study.
3 Empirical setting, data and methods